A satellite in Earth orbit has a mass of 100kg and is at an altitude of 2.00 x 10^6 m.
(a) What is the potential energy of the satellite-Earth system? (b) What is the magnitude of the gravitational force exerted by the earth on the satellite?
1
Expert's answer
2021-02-07T19:17:39-0500

Answer

a) the potential energy of the satellite-Earth system

"U=-\frac{GMM'}{r}"

Where M is mass of earth and othercof satellite.

"U=-\frac{ 6.67×\times10^{−11} \times5.97 × 10^{24}}{2\times10^6 }"

"=" "-20\times10^{7} J"

b) magnitude of the gravitational force exerted by the earth on the satellite

"F=\frac{GMM'}{r^2}"

"=\frac{ 6.67×\times10^{−11} \times5.97 × 10^{24}}{(2\times10^6 )^2}"

"=100N"